OpenStudy (anonymous):

Two mechanics worked on a car. The first mechanic worked for 20 hours, and the second mechanic worked for 15 hours. Together they charged a total of $3225. What was the rate charged per hour by each mechanic if the sum of the two rates was $190 per hour? First mechanic: $ per hour Second mechanic :$ per hour Will medal

OpenStudy (anonymous):

@Vocaloid

Vocaloid (vocaloid):

first off, lets define our variables f = number of hours worked by the first mechanic s = number of hours worked by the second mechanic, with me so far?

OpenStudy (anonymous):

yes

Vocaloid (vocaloid):

actually, let me make a quick correction

Vocaloid (vocaloid):

f = the rate charged by the first mechanic s = the rate charged by the second mechanic

Vocaloid (vocaloid):

anyway, it says that the sum of the rates is $190 so f + s = 190, with me so far?

OpenStudy (anonymous):

yes

Vocaloid (vocaloid):

now, it says that they charged 3225 together, total so now, we can write another equation first mechanic money + second mechanic money = 3225

Vocaloid (vocaloid):

so, now we can write an expression to figure out how much each mechanic made the first mechanic worked 20 hours, so he charges: 20f the second mechanic worked 15 hours, so he charges: 15s

OpenStudy (anonymous):

yes

Vocaloid (vocaloid):

that gives us two equations: f + s = 190 20f + 15s = 3225 solve for f and s

OpenStudy (anonymous):

but how

Vocaloid (vocaloid):

start by multiplying the first equation by 20